OpenStudy (anonymous):

Find the sum of the arithmetic sequence. -10, -7, -4, -1, 2, 5, 8 11 -70 0 -7

OpenStudy (anonymous):

11

OpenStudy (anonymous):

thank you!! :)

OpenStudy (anonymous):

8 +3=11

OpenStudy (anonymous):

@DaWizjr i think you're going wrong. 8+5+2 = 15, -10-7-4-1 = -22 -22+15 = -7 is the answer @n14r96

OpenStudy (anonymous):

ok thank you! :)